# Ant Colony Sampling with GFlowNets for Combinatorial Optimization

### v2

@@ Abstract
This paper introduces the Generative Flow Ant Colony Sampler (GFACS), a neural-guided probabilistic search algorithm for solving combinatorial optimization (CO). GFACS integrates generative flow networks (GFlowNets), an emerging amortized inference method, with ant colony optimization (ACO), a promising probabilistic search algorithm. Specifically, we use GFlowNets to learn a constructive policy in combinatorial spaces for enhancing ACO by providing an informed prior distribution over decision variables conditioned on input graph instances. Furthermore, we introduce a novel off-policy training algorithm for scaling conditional GFlowNets into large-scale combinatorial spaces by leveraging local search and shared energy normalization. Our experimental results demonstrate that GFACS outperforms baseline ACO algorithms in seven CO tasks and is competitive with problem-specific heuristics for vehicle routing problems.
@@ 1 Introduction
Combinatorial optimization (CO) is a crucial domain involving high-impact applications, including vehicle routing, hardware design optimization, and biological sequence design. Ant colony optimization (ACO) [14, 68] is a general-purpose meta-heuristic algorithm designed to address these tasks. ACO employs a constructive solution generation scheme that sequentially adds elements to a partial solution, known for its flexibility across CO tasks. This approach involves collecting experiences from multiple agents (ants) to iteratively update a distribution (pheromone update) based on a prior distribution known as the heuristic matrix. The major benefits of ACO include: (1) its robustness against local optima, enabled by the algorithm’s stochastic nature, which dynamically balances exploration and exploitation, preventing premature convergence [14]; and (2) its flexible solution generation scheme, which allows it to effectively solve a wide range of combinatorial optimization tasks [79].
The main challenge with ACO lies in designing effective prior distributions (i.e., the heuristic matrix) for their probability models. Inadequate priors can lead to time-consuming pheromone update iterations. Recent advancements, known as DeepACO [79], address this by using policy gradient reinforcement learning (RL) [76, 71] to pretrain the priors conditioned on given problem instances. RL trains these networks to guide highly rewarded regions, allowing generalization to new problems and enhanced performance.
Motivated by DeepACO, we propose a deep learning-based prior distribution design. However, this paper tackles challenges related to the design of prior distribution through deep amortized inference rather than the policy gradient RL method. This is motivated by the following three primary limitations in DeepACO’s use of policy gradient RL. First ( ), the on-policy nature of DeepACO restricts the flexible use of a behavior policy, potentially undermining performance by limiting effective exploration of vast combinatorial spaces. Second ( ), the reward maximizing form of RL in DeepACO struggles to capture the diverse solutions necessary in CO, even though having a variety of good candidates could significantly enhance the likelihood of finding a global optimum after local searches. Last ( ), the inherent symmetry in combinatorial solutions—where different paths can lead to identical outcomes—presents a significant challenge not adequately addressed in DeepACO.
Contribution. This paper introduces a novel ACO algorithm, the Generative Flow Ant Colony Sampler (GFACS), which addresses the limitations of DeepACO by integrating the emerging amortized inference method of GFlowNets [4, 5]. Our twofold contribution can be summarized as follows. (1) We introduce a new viewpoint to ACO using amortized inference instead of policy gradient RL. (2) Additionally, we present a novel GFlowNet training algorithm for large scale combinatorial tasks, applicable to a range of further uses over ACO.
  • ACO as off-policy amortized inference with GFlowNets. There are three reasons that GFlowNets can improve over the existing DeepACO algorithm. First ( ), the off-policy training property of GFlowNets allows for the utilization of samples drawn from any behavior policy, including powerful heuristic search algorithms commonly used in CO problems. Secondly ( ), unlike maximum reward-seeking algorithms (e.g., RL training in DeepACO), GFlowNets are sampling algorithms that are well-suited for modeling the multimodal nature of certain distributions. Lastly ( ), GFlowNets model solutions as a generative process on a directed acyclic graph (DAG), which can represent multiple trajectories that share identical states. This approach allows for accurate probabilistic modeling by considering the symmetry of combinatorial solutions [5].
  • Scaling conditional GFlowNets into large-scale combinatorial space. GFlowNets often struggle with limited scalability due to challenging exploration problems [58, 39], credit assignment issues [59, 26], and unstable training when inferring conditional distributions [82]. In combinatorial optimization, both limitations must be addressed because we deal with (1) large-scale combinatorial spaces and (2) conditional inference on problem instances. This study introduces a new off-policy training method for GFlowNets in large combinatorial spaces by effectively leveraging local search operators. It also proposes a novel training method to address instability in training conditional GFlowNets by properly normalizing conditional reward values with sample mean.
In comparative experiments, our GFACS exhibits significant superiority over the current DeepACO and classical ACO algorithms for various representative CO tasks – vehicle routing problems, scheduling problems, and a grouping problem. Moreover, GFACS outperforms or shows similar performance to competitive unsupervised learning-based vehicle routing solvers, which are designed to solve traveling salesman problems (TSP) and capacitated vehicle routing problems (CVRP) in the context of large-scale instances, such as $N=500,1000$. Lastly, we compared our approach with previous GFlowNet methods for solving combinatorial optimization problems [82, 84] in vehicle routing tasks, which are known to pose challenges for credit assignment due to their locally undecomposable properties, as highlighted by Ahn et al. [1].
2 Related works
2.1 Ant colony optimization
Ant Colony Optimization (ACO) is a meta-heuristic inspired by ant behavior in finding optimal paths. Traditional ACO algorithms, such as Ant System (AS) [14] and MAX-MIN AS [68], rely on specific heuristics for success. To reduce the reliance on heuristic design, Ye et al. [79] introduce a neural-guided ACO, called DeepACO, utilizing graph representation learning and reinforcement learning to replace manual heuristics with automatically acquired knowledge.
2.2 Neural combinatorial optimization
Neural combinatorial optimization [3] is a deep learning-based method to train an instance-conditional policy using supervised learning [75, 20, 70] or unsupervised learning [30, 41, 43, 54]. Usually, supervised learning methods are beneficial when the target problem has an accurate but time-consuming expert solver. The oracle-labeled data is collected from the oracle solvers, and deep neural networks amortize the search process, converting an expensive run-time optimization into an also expensive but one-time cost of training the neural net, allowing for scalability and faster run-time inferences than the expert solver. On the other hand, unsupervised learning proceeds through self-guided trial and error, using the COP objective as a training signal. This makes it applicable even if encountering less-studied problems where powerful expert solvers are unavailable [3, 1]. Our method can be categorized as an unsupervised learning method, which does not need expert labels. See Appendix F for a more comprehensive survey.
2.3 GFlowNets and combinatorial optimization
GFlowNets are designed for sampling combinatorial objects proportionally to a given reward function [4]. Theoretical studies have connected GFlowNets with variational inference [53] and maximum entropy reinforcement learning [56, 13]. Recent algorithmic advancements in GFlowNets have focused on developing new training objectives [52, 51, 85], enhancing credit assignment [59, 26], and improving exploration strategies [58, 63, 59, 39, 38].
GFlowNets have also found successful applications in some combinatorial optimization problems (COPs) [82, 84]. Zhang et al. [84] effectively address the symmetric nature of graph constraint optimization, demonstrating remarkable performance in tasks such as finding maximum independent sets. Despite their success, there are limitations in expanding to other combinatorial tasks due to their reliance on local credits for the learning objective proposed by Pan et al. [59]. Conversely, Zhang et al. [82] address proxy-based scheduling problems by developing a problem-specific neural architecture and introducing novel training methods for general-purpose conditional GFlowNets. One of their significant contributions is the proposal of a loss function based on the sample variance of the log partition function, known as VarGrad, which enhances the training process and performance.
@@ 3.1 Traveling salesman problem
The traveling salesman problem (TSP) is a fundamental combinatorial optimization problem, which serves as a building block for more complex CO problems like the capacitated vehicle routing problem (CVRP) and orienteering problem (OP). In light of this and for simplicity, this paper presents formulations using the two-dimensional Euclidean TSP as an illustrative example.
A TSP instance is defined on a (fully connected) graph $\mathcal{G}=\{\mathcal{V},\mathcal{D}\}$, where $\mathcal{V}$ and $\mathcal{D}$ are a node set and edge set, respectively. Each node $i\in\mathcal{V}$ has two-dimensional coordinates $v_{i}\in[0,1]^{2}$ as a feature, and each edge $(i,j)\in\mathcal{D}$ has a feature $d_{i,j}$ which is the distance between node $i$ and $j$, i.e., $d_{i,j}=\|v_{i}-v_{j}\|_{2}$. A solution to TSP, $x(\mathcal{G})$ corresponds to a Hamiltonian cycle for a given problem instance $\mathcal{G}$, which can be represented as a permutation of nodes $\pi=(\pi_{1},\ldots,\pi_{N})$. The tour length is computed as $\mathcal{E}(x(\mathcal{G}))=d_{\pi_{N},\pi_{1}}+\sum_{t=1}^{N-1}d_{\pi_{t},\pi_{t+1}}$. The optimization objective for TSP is to find $x(\mathcal{G})$ that minimizes $\mathcal{E}(x(\mathcal{G}))$.
@@ 3.2 Ant colony optimization
Ant colony optimization is a meta-heuristic to refine the solution sampling probability by an iterative procedure of solution construction and pheromone update. Given a problem instance $\mathcal{G}$, a corresponding heuristic matrix $\eta(\mathcal{G})\in\mathbb{R}^{N\times N}$ and a pheromone trails $\rho\in\mathbb{R}^{N\times N}$, a solution is constructed by auto-regressively sampling edges from the probability distribution:
@@
At each iteration, $K$ artificial ants independently sample solutions $\{\pi^{k}\}_{k=1}^{K}$ by following Equation 1. The pheromone trails $\rho$ are then adjusted using the solution set and corresponding costs, thereby updating the solution construction probability. The pheromone update rules vary depending on the ACO algorithm used. For example, in Ant System (AS) [14], all $K$ solutions contribute to the pheromone update:
@@
where $\gamma$ is the decaying coefficient (or evaporation rate) and $\Delta^{k}_{i,j}$ is the amount of pheromone laid by the $k$-th ant on the edge $(i,j)$. Here, $\Delta^{k}_{i,j}$ is proportional to the inverse of tour length if $(i,j)$ is included in $\pi^{k}$, and zero otherwise.
$\Delta^{k}_{i,j}=\begin{cases}C/f(\pi^{k})&\text{if $(i,j)\in\pi^{k}$}\\ 0&\text{otherwise}\end{cases}$ (3)
where $C$ is a constant. The smaller the cost $f(\pi^{k})$, the more pheromone is laid on the edges of $\pi^{k}$.
Note that the heuristic matrix $\eta(\mathcal{G})\in\mathbb{R}^{N\times N}$ is not updated during the search process; it only serves as a prior. ACO aligns the search process more closely with the most efficacious paths discovered; see Dorigo et al. [15] for details. In DeepACO [79], the learned graph neural network (GNN), parameterized by $\theta$, is utilized to construct the heuristic matrix $\eta(\mathcal{G},\theta)$ for a given problem instance $\mathcal{G}$, thereby providing a good prior for guiding probabilistic solution search process of ACO.
3.3 Generative flow networks
Generative Flow Networks, or GFlowNets [4, 5] provide an effective training framework for learning a policy that samples compositional objects denoted as $x\in\mathcal{X}$ towards the objective of generating $x$ with probability proportional to a given non-negative reward function $R(x)$. GFlowNets follow a constructive generative process, using discrete actions to iteratively modify a state which represents a partially constructed object. All the ways to construct these objects can be described by a directed acyclic graph (DAG), $\mathcal{H}=(\mathcal{S},\mathcal{A})$, where $\mathcal{S}$ is a finite set of all possible states, and $\mathcal{A}$ is a subset of $\mathcal{S}\times\mathcal{S}$, representing directed edges. Within this framework, we define the children of state $s\in\mathcal{S}$ as the set of states connected by edges whose head is $s$, and the parents of state $s$ as the set of states connected by edges whose tail is $s$.
We define a complete trajectory $\tau=(s_{0}\rightarrow\ldots\rightarrow s_{N})\in\mathcal{T}$ from the initial state $s_{0}$ to terminal state $s_{n}$. We denote $\tau_{{}_{\rightarrow}x}$ when the terminal state $s_{n}$ is finalized to the corresponding object $x\in\mathcal{X}$. We define the forward policy to model the forward transition probability $P_{F}(s^{\prime}|s)$ from $s$ to its child $s^{\prime}$. Similarly, we also consider the backward policy $P_{B}(s|s^{\prime})$ for the backward transition $s^{\prime}\dashrightarrow s$, where $s$ is a parent of $s^{\prime}$. We can use the forward policy to compute in a forward way the probability of a trajectory:$P_{F}(\tau;\theta)=\prod_{t=1}^{n}P_{F}\left(s_{t}|s_{t-1};\theta\right)$.
Similarly, we can use the backward policy to compute in a backward way the probability of trajectory ending at a given object $x=s_{n}$: $P_{B}(\tau_{\rightarrow x}|x;\theta)=\prod_{t=1}^{n}P_{B}\left(s_{t-1}|s_{t};\theta\right)$.
Trajectory balance [52]. Trajectory balance is the most widely used training objective for GFlowNets. When the trajectory balance objective is minimized, we achieve that the probability of sampling $x$ with the forward policy is $P_{F}^{\top}(x)\propto R(x)$, as desired, where $P_{F}^{\top}(x)=\sum_{\tau_{\rightarrow x}\in\mathcal{T}}P_{F}(\tau)$. The trajectory balance loss $\mathcal{L}_{\text{TB}}$ works by training three models, a learnable scalar $Z_{\theta}$ for the partition function, a forward policy $P_{F}(\tau;\theta)$, and a backward policy $P_{B}(\tau_{\rightarrow x}|x;\theta)$ to minimize the following objective:
$\mathcal{L}_{\text{TB}}(\tau;\theta)=\left(\log\frac{Z_{\theta}P_{F}(\tau;\theta)}{R(x)P_{B}(\tau_{\rightarrow x}|x;\theta)}\right)^{2}.$ (4)
4 Generative flow ant colony sampler (GFACS)
In this section, we provide the inference procedure for the Generative Flow Ant Colony Sampler (GFACS). GFACS is a sampling algorithm that enhances Ant Colony Optimization (ACO) by integrating it with GFlowNets. The central concept is to use GFlowNets to generate an appropriate instance-conditioned prior (edge) distribution (heuristic matrix), which more effectively guides the probabilistic search procedure of ACO.

### v4

@@ Abstract
We present the Generative Flow Ant Colony Sampler (GFACS), a novel meta-heuristic method that hierarchically combines amortized inference and parallel stochastic search. Our method first leverages Generative Flow Networks (GFlowNets) to amortize a multi-modal prior distribution over combinatorial solution space that encompasses both high-reward and diversified solutions. This prior is iteratively updated via parallel stochastic search in the spirit of Ant Colony Optimization (ACO), leading to the posterior distribution that generates near-optimal solutions. Extensive experiments across seven combinatorial optimization problems demonstrate GFACS’s promising performances.
@@ 1 Introduction
Combinatorial optimization (CO) aims to optimize an objective function over discrete, finite combinatorial structures under given constraints. It has broad applications in operations research (Zhang et al.,, 2019), scientific discovery (Naseri and Koffas,, 2020) and machine learning (Chickering et al.,, 1995, 2004). Traditionally, integer programming approaches have been extensively studied to solve CO problems (Miller et al.,, 1960). While these methods offer guaranteed optimality, they often become impractical for large-scale instances due to their exponential complexity. An alternative is to design problem-specific heuristics, but it requires domain knowledge and labor-intensive design, limiting flexibility to variant problems.
* Equal contribution. $\dagger$ Work done while Hyeonah Kim was at KAIST. Code available at: https://github.com/ai4co/gfacs. Correspondence to: Minsu Kim (minsu.kim@mila.quebec) and Sanghyeok Choi (sanghyeok.choi@kaist.ac.kr).
Recent advancements in deep learning (LeCun et al.,, 2015) brought alternative directions to CO. For a more detailed review, refer to our related works in Appendix F and the survey by Bengio et al., 2021b . Among deep learning for CO methods, supervised learning (SL) approaches distill mathematical programming methods or heuristics into powerful deep neural networks, enabling fast and scalable solution generation (Vinyals et al.,, 2015; Joshi et al.,, 2021; Sun and Yang,, 2023; Luo et al.,, 2023). However, SL methods rely on expert-labeled data, limiting their applicability as general-purpose CO solvers.
Reinforcement learning (RL) is another popular approach to CO (Khalil et al.,, 2017; Kool et al.,, 2019; Kwon et al.,, 2020; Ye et al.,, 2023). Unlike supervised learning, RL methods train deep neural networks directly using the objective function of a CO problem as a reward, without requiring expert-labeled data. This makes RL methods particularly useful because they can adapt flexibly to a wide range of CO problems, even where specific heuristics are unavailable or unexplored. However, RL requires extensive exploration of vast solution space and often involves non-trivial credit assignments across multiple decisions, which introduces significant complexity to the learning process compared to SL (Minsky,, 1961).
Hierarchical methods that leverage RL in a two-stage problem-solving procedure tackle such exploration and credit assignment problems through hierarchical decomposition. Previous works by Bello et al., (2016), Hottung et al., (2022), and Son et al., (2023) pretrain large models on massive problem instances using RL and fine-tune these models on specific problem instances at test time using a method called active search (AS). Ye et al., (2023) pretrain relatively small models with significantly reduced training times. They pretrain a neural network with RL to produce a prior distribution over solution space and iteratively update the distribution using ant colony optimization (ACO), an effective parallel stochastic search algorithm.
Contribution. In this work, we propose a novel probabilistic meta-heuristic called Generative Flow Ant Colony Sampler (GFACS), which leverages the synergy between a multi-modal amortized prior and iterative update of it using parallel stochastic search. Unlike previous approaches that use RL pretraining aimed at maximizing expected rewards, we employ Generative Flow Networks (GFlowNets; Bengio et al., 2021a, ; Bengio et al.,, 2023), designed to learn a distribution where the probability of a sample is proportional to its reward, i.e., $p(x)\propto R(x)$. This leads to a multi-modal prior distribution of solutions, allowing it to benefit more from subsequent iterative posterior updates compared to the unimodal priors pretrained with reward maximizing RL, as illustrated in Figure 1. Moreover, to enhance and stabilize GFlowNet training, we introduce several new techniques that have been verified to be significantly more effective than existing GFlowNet training methods.
Our posterior update rules follow the most progressive methods of ant colony optimization (ACO) (Ye et al.,, 2023), creating synergies between our multi-modal amortized prior and their iterative refinement via parallel stochastic search. Additionally, we also validate our method with another posterior search method, active search (AS; Hottung et al.,, 2022), demonstrating the flexibility of our approach.
Figure 1: Synergy of multi-modal prior distribution of solutions and iterative posterior update with parallel stochastic search in combinatorial optimization.
Through extensive experiments, we demonstrate the superiority of our training methods over general-purpose reinforcement learning algorithms for providing effective prior distribution on ACO search across seven combinatorial optimization benchmarks:
  • Routing problems:
    • Traveling Salesman Problem (TSP)
    • Capacitated Vehicle Routing Problem (CVRP)
    • CVRP with Time Windows (CVRPTW)
    • Prize-Collecting TSP (PCTSP)
    • Orienteering Problem (OP)
  • Scheduling problem: Single Machine Total Weighted Tardiness Problem (SMTWTP)
  • Grouping problem: Bin Packing Problem (BPP)
For problems such as the Traveling Salesman Problem (TSP) and the Capacitated Vehicle Routing Problem (CVRP), we compare our GFACS method against problem-specific reinforcement learning algorithms that utilize significantly larger and specialized deep networks.
@@ 2.1 Traveling salesman problem
The traveling salesman problem (TSP) is a fundamental combinatorial optimization problem, which serves as a building block for more complex CO problems like the capacitated vehicle routing problem and orienteering problem. In light of this and for simplicity, this paper presents formulations using the two-dimensional Euclidean TSP as an illustrative example.
A TSP instance is defined on a (fully connected) graph $\mathcal{G}=(\mathcal{V},\mathcal{D})$, where $\mathcal{V}$ and $\mathcal{D}$ are a set of nodes and edges, respectively. Each node $i\in\mathcal{V}$ has two-dimensional coordinates $v_{i}\in[0,1]^{2}$ as a feature, and each edge $(i,j)\in\mathcal{D}$ has a feature $d_{i,j}$ which is the distance between node $i$ and $j$, i.e., $d_{i,j}=\|v_{i}-v_{j}\|_{2}$. A solution to TSP, $x$ corresponds to a Hamiltonian cycle for a given problem instance $\mathcal{G}$, which can be represented as a permutation of nodes $\pi=(\pi_{1},\ldots,\pi_{N})$. The tour length is computed as $\mathcal{E}(x;\mathcal{G})=d_{\pi_{N},\pi_{1}}+\sum_{t=1}^{N-1}d_{\pi_{t},\pi_{t+1}}$. The optimization objective for TSP is to find $x$ that minimizes $\mathcal{E}(x;\mathcal{G})$.
2.2 Generative flow networks
Generative Flow Networks, or GFlowNets (Bengio et al., 2021a, ; Bengio et al.,, 2023), provide an effective training framework for learning a policy that samples compositional objects $x\in\mathcal{X}$. The objective is to generate $x$ with probability proportional to a given non-negative reward function $R(x)$. GFlowNets follow a constructive generative process, using discrete actions to iteratively modify a state which represents a partially constructed object. All the ways to construct these objects can be described by a directed acyclic graph (DAG), $\mathcal{H}=(\mathcal{S},\mathcal{A})$, where $\mathcal{S}$ is a finite set of all possible states, and $\mathcal{A}$ is a subset of $\mathcal{S}\times\mathcal{S}$, representing directed edges. Within this framework, we define the children of state $s\in\mathcal{S}$ as the set of states connected by edges whose head is $s$, and the parents of state $s$ as the set of states connected by edges whose tail is $s$.
We define a (complete) trajectory $\tau=(s_{0}\rightarrow\ldots\rightarrow s_{N})\in\mathcal{T}$ from the initial state $s_{0}$ to terminal state $s_{n}$. We define the forward policy to model the forward transition probability $P_{F}(s^{\prime}|s)$ from $s$ to its child $s^{\prime}$. Similarly, we also consider the backward policy $P_{B}(s|s^{\prime})$ for the backward transition $s^{\prime}\dashrightarrow s$, where $s$ is a parent of $s^{\prime}$. We can use the forward policy to compute in a forward way the probability of a trajectory: $P_{F}(\tau;\theta)=\prod_{t=1}^{n}P_{F}\left(s_{t}|s_{t-1};\theta\right)$. Similarly, we can use the backward policy to compute in a backward way the probability of trajectory ending at a given object $x=s_{n}$: $P_{B}(\tau|x;\theta)=\prod_{t=1}^{n}P_{B}\left(s_{t-1}|s_{t};\theta\right)$.
Trajectory balance (Malkin et al.,, 2022).
Trajectory balance (TB) is the most widely used training objective for GFlowNets. It works with three key components: a learnable scalar $Z_{\theta}$ for the partition function, a learnable forward policy $P_{F}(\tau;\theta)$, and a backward policy $P_{B}(\tau|x)$ that can either be learnable or fixed with a predefined configuration (Zhang et al., 2022a, ). TB loss to minimize for any trajectory $\tau$ is defined as:
$\mathcal{L}_{\text{TB}}(\tau;\theta)=\left(\log\frac{Z_{\theta}P_{F}(\tau;\theta)}{R(x)P_{B}(\tau|x)}\right)^{2}.$ (1)
When the trajectory balance loss is minimized to zero for all complete trajectories, the probability of sampling $x$ with the forward policy is $P_{F}^{\top}(x)\propto R(x)$, as desired, where $P_{F}^{\top}(x)=\sum_{(\tau\rightarrow x)\in\mathcal{T}}P_{F}(\tau)$ with $(\tau\rightarrow x)$ indicating that the terminal state of $\tau$ is $x$.
@@ 2.3 Ant colony optimization
Ant colony optimization (ACO; Dorigo et al.,, 2006) is a meta-heuristic to refine the distribution over solution space by an iterative procedure of solution construction and pheromone update. Given a problem instance $\mathcal{G}$, a corresponding heuristic matrix $\eta(\mathcal{G})\in\mathbb{R}^{N\times N}$ and a pheromone trails $\rho\in\mathbb{R}^{N\times N}$, a solution is constructed by auto-regressively sampling edges from the probability distribution:
@@
At each iteration, $K$ artificial ants independently sample solutions $\{\pi^{k}\}_{k=1}^{K}$ by following Eq.(2). The pheromone trails $\rho$ are then adjusted using the sampled solutions and their corresponding costs, thereby updating the solution distribution. The exact pheromone update rules depend on the specific ACO algorithm being used. As a representative example, in Ant System (Dorigo et al.,, 1996), all $K$ solutions contribute to the pheromone update as follows:
@@
where $\gamma$ is the decaying coefficient (or evaporation rate) and $\Delta^{k}_{i,j}$ is the amount of pheromone laid by the $k$-th ant on the edge $(i,j)$. The value of $\Delta^{k}_{i,j}$ is proportional to the inverse of tour length if edge $(i,j)$ is part of $\pi^{k}$, and zero otherwise:
$\Delta^{k}_{i,j}=\begin{cases}C/f(\pi^{k})&\text{if $(i,j)\in\pi^{k}$}\\ 0&\text{otherwise},\end{cases}$ (4)
where $C$ is a constant. The smaller the cost $f(\pi^{k})$, the more pheromone is laid on the edges of $\pi^{k}$. This mechanism encourages future ants to follow promising paths that lead to high-quality solutions.
Note that the heuristic matrix $\eta(\mathcal{G})$ is not updated during the search process; it only serves as an initial prior. This prior is augmented with the pheromone $\rho$ that is typically initialized as a constant matrix and then iteratively updated, resulting in the posterior solution distribution that converges to the Dirac-delta distribution of optimal solution under certain conditions (Dorigo and Gambardella,, 1997; Dorigo and Blum,, 2005).
In DeepACO (Ye et al.,, 2023), the learned graph neural network (GNN), parameterized by $\theta$, is utilized to construct the heuristic matrix $\eta(\mathcal{G},\theta)$ for a given problem instance $\mathcal{G}$, providing an informative prior for guiding probabilistic solution search process of ACO. The limitation of the approach is that they use REINFORCE (Williams,, 1992) for GNN training, which suffers from mode-collapse, restricting further improvement from the subsequent iterative pheromone update. Our work seeks to overcome this limitation by incorporating a diversified, multi-modal solution distribution, which shares similar motivation with approximate global optimization methods such as simulated annealing (Kirkpatrick et al.,, 1983).
3 Generative flow ant colony sampler
In this section, we provide the inference procedure for the Generative Flow Ant Colony Sampler (GFACS). GFACS enhances Ant Colony Optimization (ACO) by integrating it with multi-modal prior from GFlowNet-based training.